OpenStudy (anonymous):

Write an equation in standard form for a line with an m = -1/2 and the point (-2, -3).

OpenStudy (anonymous):

Ok so the first think you have to do is get an equation of some sort... i recommend slope intercept... for now. You will convert to standard form later. Plug in your x and y values into y=mx+b and plug in your m and solve for b (your y-intercept). Once you get your slope intercept form, you will convert to standard form. Here is a link on how to do that... there are some practice problems to work with too. I hope this helps!!! http://www.mathwarehouse.com/algebra/linear_equation/slope-intercept-to-standard-form.php

OpenStudy (anonymous):

Or you could put this in point slope form. \[y - y _{1} = m(x - x _{1})\]

OpenStudy (anonymous):

m = -1/2 y1 = -3 x1 = -2

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.